OpenStudy (anonymous):

Okay i need 2 even Functions and 2 Odd Function and they have to be complex. anyone got one. P.S I have to prove that they are even and odd functions and i having difficulty finding a complex function

OpenStudy (anonymous):

Even function f(x) = f(-x) Odd function f(x) = -f(-x) And they have to be complex functions also, which means they have to include complex numbers. We can think about some examples: (1) Even functions \[y = i.x^2\] \[y = i.x^4\] (2) Odd functions \[y = i.\sin(x)\] \[y = i.\tan(x)\]

OpenStudy (anonymous):

You can prove them by the above definition.

OpenStudy (anonymous):

oh w8 maybe i didnt make the question clear, when i meant my complex i ment by hard lol

OpenStudy (anonymous):

im only in grade 12 never learned imaginary numbers

OpenStudy (anonymous):

Ok how hard you need :) lol Is it even function hard enough for you? \[y = 5x^6 -7x^4 + 100x^2-10\]

OpenStudy (anonymous):

yup, do u have another one and i also need 2 odd functions

OpenStudy (anonymous):

Ok, just the basic idea is x^2 is even function and sin(x) is odd function examples. So from 2 basics, I can develop any kind of thinking. The following is an even function, right? \[y = 0.5\tan(x) - 100\sin(x) + x^3\] Try to develop more by yourself. I think with 2 basics idea, it is not difficult for you, a 12-grade student.

OpenStudy (anonymous):

oh ok thanks

OpenStudy (anonymous):

I am sorry, the above answer with tan(x) is odd function (my mistake) :P

OpenStudy (anonymous):

alright thanks

OpenStudy (anonymous):

wait would y=cos(x) be a even function cuz it symmetrical about the y-axis?

OpenStudy (anonymous):

About cos(x). You can easily see that: cos(-x) = cos(x). So cos(x) is an even function. But because it is symmetrical through x-axis, not y-axis